OpenStudy (anonymous):

The sum of three numbers is . The second number is times the third. The third number is less than the first . What are the numbers?

OpenStudy (chihiroasleaf):

I think your question is not complete.. :)

OpenStudy (anonymous):

oops! sorry haha here The sum of three numbers is 96 . The second number is3 times the third. The third number is 6 less than the first . What are the numbers?

OpenStudy (chihiroasleaf):

let the three numbers are x,y,z from the first sentence, "The sum of three numbers is 96" how do you say this in x,y,z ?

OpenStudy (anonymous):

not sure i understand what you mean by how we say this in x,y,z

OpenStudy (chihiroasleaf):

let's say that the three numbers are x,y, and z. the first sentence say that "The sum of three numbers is 96" , we can write the sentence as "x+y+z = 96" it's what I mean.. :) do you understand now?

OpenStudy (anonymous):

oh ok yesss sorry haha so x,y, and z would be 32 bcause 96/3 =32

OpenStudy (chihiroasleaf):

wait..., we have no information that x, y, and z are equals... from the second sentence, we have " The second number is3 times the third" if y is the second number, and z is the third number, how do you write this in math? (in y and z)

OpenStudy (anonymous):

3z=y ?

OpenStudy (chihiroasleaf):

correct... :) now, from the third sentence, we have " The third number is 6 less than the first " let, x be the first number and z be the third number, how do you write it in term of x and z?

OpenStudy (anonymous):

x-6=z

OpenStudy (chihiroasleaf):

okay.., x - 6 = z, so x = z + 6, do you get it so far?

OpenStudy (chihiroasleaf):

now, we have three equations x+y+z = 96 3z = y x = z + 6 now, substitute (change) the value of x and y in the first equation with the second and third equation, what do you get?

OpenStudy (anonymous):

yes =) however i feel like that was the easy part.. and the hard part is coming =o